OpenStudy (anonymous):

Hi, I have a summer math packet that is focusing on discrete math and is a review of precalculus and algebra at a high school level. Does anyone know of any good websites I could use for help? Thanks.

OpenStudy (anonymous):

Well, This one. if there's a concept or even just a single equation you don't understand, post it here and get a good descriptive answer pretty quick.

OpenStudy (anonymous):

Okay here's one. I have a pretty simple equation to simplify; \[2k/2k+4=k/x+r\] I feel like this should be simple but my brain is fried from doing no math over the summer, unfortunately.

OpenStudy (anonymous):

Correction: I have to identify if they are true statements are not. Sorry.

OpenStudy (anonymous):

wolframalpha.com is perfect for precalc and algebra

OpenStudy (anonymous):

"2k/2k+4=k/x+r" becomes "1+4=k/x+r" then "5=k/x+r", being that this equation has 3 unknowns and i assume the domain is integers, then the statement is true since you can find at least one number for k,x and r to make the statement true for example, k=10,x=5, r=3 "5=10/5+3"

OpenStudy (anonymous):

patrickjmt.com. Dude does really awesome videos on pretty much anything you will need.

OpenStudy (anonymous):

Thanks xon for the reply, but I made an error. instead of r, it should be 4. My bad. ><;